# Belo Horizonte, MG, Brazil

### v3

@@ 5.1 Areal Data: Respiratory Disease Hospitalization
We first analyze a dataset on the count of hospital admissions due to respiratory disease in 2010 for the 134 intermediate zones (IZ) to the north of the river Clyde in the Great Glasgow and Clyde health board in Scotland. This dataset is publicly available in theCARBayesdata package (Lee,, 2022). For each IZ, the expected number of admissions based on the national age- and sex-standardized rates, alongside the percentage of people classified as income deprived in each IZ are also available. The standardized morbidity ratio (SMR), the ratio between the number of observed and expected cases, is often used to control for confounding factors when estimating risks (Greenland,, 1984). The goals of the analyses are: (1) to provide smooth estimates of the SMR, adjusted for the percentage of income deprivation, and (2) to contrast the outcomes from the proposed methodology with other established methods.
@@
In this application, $Y(\mathbf{s}_{i})$ is the number of admissions by respiratory disease registered at the $i$-th IZ, denoted $\mathbf{s}_{i}$. In the $i$-th IZ, define $x_{i}$ as the percentage of people who are defined to be income deprived and $E_{i}$ as the expected number of admissions. We analyze the data using the GLMM model from Section 3.1, with a Poisson likelihood, a $\log$ link function, an intercept, a covariate $x_{i}$, and an offset $E_{i}$. We employed the homoscedastic HGP (with a PEC function with smoothness $\nu=0.7$) as the prior for the random effects; the choice of $\nu=0.7$ is justified in Table S.4 of the Appendix Material. The priors for the parameters in the HGP GLMM were set to be those presented in Section 3.2. Inference was performed using four parallel chains with 5000 MCMC samples, thinned by 10, after discarding the initial 1000 iterations as the warm-up period. The initialization of the algorithm is described in Section 3.2. The highest split-$\hat{R}$ was smaller than $1.01$, providing evidence of convergence.
@@
Two competing models, BYM and DAGAR, were considered as the priors for the spatial random effects $\mathbf{z}$ for comparison, as outlined in Sections 4.1 and S.2.1, respectively. In the BYM model, parametrized by the marginal precision (Riebler et al.,, 2016), we followed Morris et al., (2019) and set standard normal priors for the intercept ($\beta_{0}$), regression coefficient ($\beta_{1}$), and the $\mathrm{logit}$ of the mixing parameter $\xi$. Similarly, a standard normal distribution truncated on positive real numbers was used for $\sigma$. The priors for the DAGAR model are the ones presented in Section 4.1. The initialization of the MCMC algorithm followed the same principle used for the HGP model. In both cases, we used four parallel chains with a warm-up period of 4000 iterations and based the inferences on 20000 MCMC samples, thinned by 20. For these models, the parameters split-$\hat{R}$ provided evidence of convergence. In addition, trace and density plots for posterior samples of all fitted model parameters are provided in Section S.5.1 of the Appendix Material.
@@
We present the posterior medians and 95% CIs for the parameters from the HGP, DAGAR, and BYM models in Table 2. When considering the coefficient linked to the covariate representing income deprivation, all three models yield similar posterior estimates. The positive posterior estimates the respective parameter ($\beta_{1}$) indicate a positive association between proportion of income deprived people and hospitalization rates for respiratory diseases. Among these models, DAGAR exhibits the broadest credible intervals for the intercept, whereas the marginal posterior distribution for this parameter in the BYM model is highly concentrated around 0.22-0.22. Examining the spatial aspect, the confidence interval for the practical range parameter $\rho$ in the HGP model implies that the distance at which the spatial correlation between random effects from two distinct spatial units decreases to $0.10$ is at most $6.95$ kilometers. Conversely, the 95% CI for the spatial dependence parameter $\psi$ from the DAGAR model spans from $0.07$ to $0.83$. In other words, this credible interval suggests that the clustering effect ranges from almost negligible to very strong. A similar pattern is observed for the parameter $\xi$ in the BYM model.
Figure 5: Density histogram of the pairwise Hausdorff distances observed in the dataset and. In blue, estimated posterior density of the practical range $\rho$, and, in red, the prior used for $\rho$.
To gain a better understanding of this value in the context of the data, Figure 5 displays a density histogram of all pairwise Hausdorff distances observed in the dataset. The plot also features the prior distribution (in red) and estimated posterior density (in blue) associated with the practical range. Notably, the majority of the posterior distribution mass for $\rho$ falls below the concentration of most pairwise distances, indicating a negligible clustering effect in hospital admissions. An additional analysis of the spatial correlation is presented in Section S.5 in the Appendix Material.
In Table 2, we also present two model comparison criteria described in Section 4 for the models fitted to the data. These results suggest the HGP stands out as the best fit based on the LOOIC, rendering the HGP prior a competitive fit for these data. Despite the modest improvement in model fitting capacity, the HGP possesses an advantage in terms of intuitive interpretability of its spatial dependence, achieved in the last two paragraphs through the practical range parameter $\rho$. Among the specialized models for areal data, the BYM was the least adequate in this application. However, none of the models delivered a remarkably better fit to this dataset. This is further reinforced by looking at the maps of observed and smoothed SMR estimates presented in Figure S.3 from the Appendix Material.
@@ 5.2 Data fusion: Air pollution in Ventura and Los Angeles
Figure 6: PM2.5 in Ventura and Los Angeles counties in California, United States. Points represent direct measurements. Tiles are satellite derived estimates.
Our second application analyzes fused fine particulate matter (PM2.5) data in Ventura and Los Angeles counties, California, USA. The dataset combines point-referenced direct measurements (U.S. Environmental Protection Agency Clean Air Markets Division,, 2018) from 1999 to present with areal satellite-derived estimates (Van Donkelaar et al.,, 2015), averaged from 2010 to 2012 (Figure 6). To align temporally, we averaged the direct measurements over the same period. The study region comprises 19 measurement stations and 184 grid boxes (approximately 101.95 km2 each), excluding islands. Our primary goal is to predict PM2.5 concentrations at a finer scale (approximately 45 km2) using both data sources. Secondarily, we aim to estimate average PM2.5 concentration and its spatial dependence across both counties.
Building on previous concepts, we introduce the notation for this application. The response variable is the recorded PM2.5, noted $Y(\mathbf{s}_{i})$ at the spatial unit $\mathbf{s}_{i}$, which is either a point or a grid box. We again resort to the GLMM model from Section 3.1. In this case, there are no covariates, the likelihood is Normal with variance $\tau^{2}$, and the link function is the identity. A heteroscedastic HGP prior is placed on the random effect allowing for differing variances between point-referenced and areal data through the SD function from Equation (2) with $w(\mathbf{s})=\mathbbm{1}(\mathcal{A}(\mathbf{s})>0)$. For interpretability, we present results for $\sigma=\exp\{\alpha_{0}\}$ and $\sigma_{a}=\exp\{\alpha_{0}+\alpha_{1}\}$ instead of $\alpha_{0}$ and $\alpha_{1}$. The model was fit with $\nu$ fixed at different values, and we selected the model with best goodness-of-fit ($\nu=0.8$). For more details, see Table S.5 in the Appendix Material. The priors for model parameters are as in Section 4.2. The MCMC algorithm initialization is described in Section 3.2. We used four parallel chains, each comprising a warm-up period of 2000 samples and an additional 2000 samples for inference. All model parameters exhibited split-$\hat{R}$ values below, indicating good convergence. This is further supported by the trace plots presented in Section S.6.1 of the Appendix Material.
For comparison, we also fit an AGP random effect model, with priors for the model parameters as described in Section 4.2. Two meshes of different resolutions were considered; see Figure S.17 in the Appendix Material. The sparse- and fine-mesh AGP models are denoted AGP1 and AGP2, respectively.
Table 3: Parameter estimates and out-of-sample prediction assessment for the three models fitted to the fused dataset. Marginal posterior summary through median and 95% credible intervals. RMSP is the root mean squared error of prediction; CPP is the frequentist coverage of the 95% prediction interval; Width is the width of the prediction interval; and IS is the interval score.
@@
$\beta$ 5.605 (4.688, 6.449) 6.219 (2.16, 10.10) 6.187 (5.879, 6.478)
$\rho$ 13.826 (7.817, 23.608) 13.163 (5.140, 30.244) 0.632 (0.456, 0.832)
$\tau$ 0.178 (0.073, 0.312) 1.394 (1.242, 1.563) 0.539 (0.394, 0.719)
$\sigma$ 3.849 (2.921, 5.181) 1.703 (0.960, 2.910) 2.399 (2.024, 2.837)
$\sigma_{a}$ 1.241 (1.037, 1.505)
RMSP 1.05 1.45 1.64
Width 3.57 2.53 9.67
CPP 95.5 78.6 95.5
IS 4.80 15.11 13.65
We present the posterior median and 95% CI of the model parameters in Table 3. When looking at the common parameters between the three models (namely, $\beta_{0}$, $\sigma$, $\rho$, and $\tau$), the HGP acts as a compromise between AGP1 and AGP2 while eliminating the arbitrary choice of the level of discretization for approximating numerical integrals. Unlike the aggregated models, our method allows the data to tell us whether the variance of the underlying spatial parameter is smaller or larger than the same quantity for point-referenced data. Interestingly, the estimated SD of the point-referenced data is 3.09 times larger then the areal data (95% CI: 2.43, 4.12). The interpretation of $\rho$, presented in tens of kilometers, in both models is carried out in a similar fashion because the Hausdorff distance between points is equivalent to the distance between points used by the AGP models.
Table 3 also presents the out-of-sample prediction performance of the three models, assessed via 10-fold cross-validation. Using four metrics from our prior simulation study (RMSP, CPP, prediction interval width, IS), we find the HGP offers the most reliable point predictions, evidenced by its lowest RMSP. The HGP also maintains nominal coverage for its prediction intervals. While the AGP2 achieves similar coverage, its prediction intervals are considerably wider, resulting in inferior interval scores. Table S.6 in the Appendix Material provides a breakdown by data type, highlighting the poor performance of competing methods for point-referenced data. The PI frequentist coverage, for example, can be as low as 3%.
The drastic differences in the AGP results for different mesh resolutions are striking. The credible interval width for the intercept varies significantly between the two models. Crucially, the practical range reveals a key divergence: AGP1 indicates moderate spatial dependence, while AGP2 suggests weak dependence, likely explaining the observed overfitting in the latter. This weak dependence may imply AGP2 assigns nearly independent random intercepts to each data point, with variances inversely proportional to the sample unit’s area. AGP mesh resolution is typically practitioner-informed, and results are highly sensitive to this tuning parameter, which lacks a consensus on “optimal resolution” (Fuentes and Raftery,, 2005; Liu et al.,, 2011; Wang and Furrer,, 2021).
@@ 6 Discussion
The proposed framework unifies the modeling of spatial data, extending geostatistical models to include areal and fused data within the same setting. As a prior for spatial random effects in a GLMM, the HGP proves to be competitive to specialized models for diverse spatial data types. Our simulation studies indicate that the HGP is particularly advantageous for areal data with units of varying sizes and shapes, as confirmed by our disease mapping application. Unlike its specialized counterparts, the HGP also provides a directly interpretable spatial parameter, offering insights into the practical range of spatial dependence. In data fusion contexts, the HGP delivers more reliable interval predictions compared to the sparse AGP, and it provided similar results to the fine AGP in parameter estimation and spatial range. Additionally, in the air pollution data application, the proposed model consistently outperformed the AGP regardless of the resolution of its mesh. This success is even more remarkable considering the HGP bypasses the subjective step of defining a grid or mesh, a crucial but problematic requirement for the AGP. In addition, independent of the spatial data type, it continues to have direct and interpretable parameters.
Our proposed methodology shares the “big nn” computational complexity challenge inherent in traditional geostatistical approaches. Consequently, the computational cost of evaluating the finite-dimensional likelihood induced by the HGP for a set of $n$ areal units scales as $\mathcal{O}(n^{3})$. Proving the validity of the PEC function under the Hausdorff distance on Euclidean space was a demanding theoretical task. Relaxing the isotropy assumption for this function presents further theoretical complexities. Moreover, to apply the HGP to data measured on manifolds (like the sphere), new theoretical results are needed to guarantee the positive definiteness of a correlation function within that specific geometric context.
The HGP framework offers numerous opportunities for development. Further researching its theoretical properties will strengthen the foundation for its applications. A crucial focus should be developing correlation functions designed specifically for the HGP, particularly those suited to scenarios where spatial data is observed on a sphere. Approximations inspired by methods like the Nearest Neighbor Gaussian Process (Datta et al.,, 2016; Quiroz et al.,, 2023) or reduced rank approximations (Solin and Särkkä,, 2020) could significantly broaden the applicability of the HGP to large-scale datasets. These advancements hold the potential to establish the HGP as an even more powerful and versatile tool for spatial data analysis.

### v4

@@ 5.1 Areal Data: Respiratory Disease Hospitalization
We first analyze a dataset on the count of hospital admissions due to respiratory disease in 2010 for the 134 intermediate zones (IZ) to the north of the river Clyde in the Great Glasgow and Clyde health board in Scotland. This dataset is publicly available in the CARBayesdata package (lee2022carbayesdata). For each IZ, the expected number of admissions based on the national age- and sex-standardized rates, alongside the percentage of people classified as income deprived in each IZ are also available. The standardized morbidity ratio (SMR), the ratio between the number of observed and expected cases, is often used to control for confounding factors when estimating risks (greenland1984bias). The goals of the analyses are: (1) to provide smooth estimates of the SMR, adjusted for the percentage of income deprivation, and (2) to contrast the outcomes from the proposed methodology with other established methods.
@@
In this application, $Y(\mathbf{s}_{i})$ is the number of admissions by respiratory disease registered at the $i$-th IZ, denoted $\mathbf{s}_{i}$. In the $i$-th IZ, define $x_{i}$ as the percentage of people who are defined to be income deprived and $E_{i}$ as the expected number of admissions. We analyze the data using the GLMM model from Section 3.1, with a Poisson likelihood, a $\log$ link function, an intercept, a covariate $x_{i}$, and an offset $E_{i}$. We employed the homoscedastic HGP (with a PEC function with smoothness $\nu=0.7$) as the prior for the random effects; the choice of $\nu=0.7$ is justified in Table S.4 of the Supplementary Material. The priors and initialization scheme for the parameters in the HGP GLMM match those presented in Section 3.2. Inference was performed using four parallel chains with 5000 MCMC samples, thinned by 10, after discarding the initial 1000 iterations as the warm-up period. The highest split-$\hat{R}$ was smaller than $1.01$, providing evidence of convergence.
@@
Two competing models, BYM and DAGAR (see the Appendix for further details), were considered as the priors for the spatial random effects $\mathbf{z}$ for comparison. In the BYM model, parametrized by the marginal precision (riebler2016intuitive), we followed morris2019Bayesian and set standard normal priors for the intercept ($\beta_{0}$), regression coefficient ($\beta_{1}$), and the $\mathrm{logit}$ of the mixing parameter $\xi$. Similarly, a standard normal distribution truncated on positive real numbers was used for $\sigma$. The priors for the DAGAR model are the ones presented in Section S.4.1. The initialization of the MCMC algorithm followed the same principle used for the HGP model. In both cases, we used four parallel chains with a warm-up period of 4000 iterations and based the inferences on 20000 MCMC samples, thinned by 20. For these models, the parameters split-$\hat{R}$ provided evidence of convergence. In addition, trace and density plots for posterior samples of all fitted model parameters are provided in Section S.6.1 of the Supplementary Material.
@@
We present the posterior medians and 95% CIs for the parameters from the HGP, DAGAR, and BYM models in Table 2. When considering the coefficient linked to the covariate representing income deprivation, all three models yield similar posterior estimates. All models found a positive association between income deprivation and hospitalizations ($\beta_{1}$). However, the HGP provided a more precise and interpretable estimate of spatial dependence via its practical range parameter, whereas the credible intervals for the dependence parameters in the BYM and DAGAR models were very wide, suggesting inconclusive evidence of spatial clustering.
To gain a better understanding of this value in the context of the data, in the Supplementary Material, Figure S.9 displays a density histogram of all pairwise Hausdorff distances observed in the dataset. The plot also features the prior distribution (in red) and estimated posterior density (in blue) associated with the practical range. Notably, the majority of the posterior distribution mass for $\rho$ falls below the concentration of most pairwise distances, indicating a negligible clustering effect in hospital admissions. An additional analysis of the spatial correlation is presented in Section S.6 in the Supplementary Material.
In Table 2, we also present two model comparison criteria described in Section 4 for the models fitted to the data. These results suggest the HGP stands out as the best fit based on the LOOIC, rendering the HGP prior a competitive fit for these data. Despite the modest improvement in model fitting capacity, the HGP possesses an advantage in terms of intuitive interpretability of its spatial dependence, achieved in the last two paragraphs through the practical range parameter $\rho$. Among the specialized models for areal data, the BYM was the least adequate in this application. However, none of the models delivered a remarkably better fit to this dataset. This is further reinforced by looking at the maps of observed and smoothed SMR estimates presented in Figure S.8 from the Supplementary Material.
@@ 5.2 Data fusion: Air pollution in Ventura and Los Angeles
Our second application analyzes fused fine particulate matter (PM2.5) data in Ventura and Los Angeles counties, California, USA. The dataset combines point-referenced direct measurements (epa2022clean) with areal satellite-derived estimates (van2015global), averaged from 2010 to 2012 (Supplementary Material, Figure S.23). To align temporally, we averaged the direct measurements over the same period. The study region comprises 19 measurement stations and 184 grid boxes (approximately 101.95 km2 each), excluding islands. Our primary goal is to predict PM2.5 concentrations at a finer scale (approximately 45 km2) using both data sources. Secondarily, we aim to estimate average PM2.5 concentration and its spatial dependence across both counties.
Building on previous concepts, we introduce the notation for this application. The response variable is the recorded PM2.5, denoted $Y(\mathbf{s}_{i})$. In this case, $\mathbf{s}_{i}$ is either a point or a grid box. We again resort to the GLMM model from Section 3.1, but without any covariates. We assume the likelihood is Normal with variance $\tau^{2}$, and the link function is the identity. A heteroscedastic HGP prior is placed on the random effect allowing for differing variances between point-referenced and areal data through the SD function from Equation (2) with $w(\mathbf{s})=\mathbbm{1}(\mathcal{A}(\mathbf{s})>0)$. For interpretability, we present results for $\sigma=\exp\{\alpha_{0}\}$ and $\sigma_{a}=\exp\{\alpha_{0}+\alpha_{1}\}$ instead of $\alpha_{0}$ and $\alpha_{1}$. The model was fit with $\nu$ fixed at different values, and we selected the one with best goodness-of-fit ($\nu=0.7$). For more details, see Table S.5 in the Supplementary Material. The priors for model parameters are as in Section 4.2. We used four parallel chains, each comprising a warm-up period of 1000 samples and an additional 1000 samples for inference. All model parameters exhibited split-$\hat{R}$ values below 1, indicating good convergence. This is further supported by the trace plots presented in Section S.7.1 of the Supplementary Material.
For comparison, we also fit two models with AGP random effects, with priors for the model parameters as described in Section 4.2. The two models only differ in the precision of the numerical integrals approximation. The sparse- and fine-resolution AGP models are denoted AGP1 and AGP2, respectively.
Table 3: Parameter estimates and out-of-sample prediction assessment for the three models fitted to the fused dataset. Marginal posterior summary through median and 95% credible intervals. RMSP is the root mean squared error of prediction; CPP is the frequentist coverage of the 95% prediction interval; and IS is the interval score.
@@
$\beta$ 5.593 (4.659, 6.431) 6.176 (2.160, 9.990) 6.344 (6.102, 6.585)
$\rho$ 15.076 (8.445, 28.561) 13.093 (4.809, 29.720) 0.493 (0.340, 0.750)
$\tau$ 0.174 (0.078; 0.324) 1.403 (1.253; 1.573) 0.539 (0.394, 0.719)
$\sigma$ 3.893 (2.962; 5.359) 1.717 (0.933; 2.938) 2.877 (2.144; 3.566)
$\sigma_{a}$ 1.275 (1.064; 1.590)
RMSP 2.22 3.55 4.45
CPP 95.6 77.3 93.6
IS 5.08 15.4 9.55
We present the posterior median and 95% CI of the model parameters in Table 3. When looking at the common parameters between the three models (namely, $\beta_{0}$, $\sigma$, $\rho$, and $\tau$), the HGP acts as a compromise between AGP1 and AGP2 while eliminating the arbitrary choice of the level of discretization for approximating numerical integrals. Unlike the aggregated models, our method allows the data to tell us whether the variance associated to areal units is smaller or larger than that of point-referenced data. Interestingly, the estimated SD of the point-referenced data is 3.09 times larger than the areal data (95% CI: 2.43, 4.12). The interpretation of $\rho$, presented in tens of kilometers, in both models is carried out in a similar fashion because the Hausdorff distance reduces to the Euclidean distance between points at the point-level.
Additionally, Table 3 presents the out-of-sample prediction performance of the three models, assessed via 10-fold cross-validation. Using the three metrics from our prior simulation study (RMSP, CPP, IS), we find the HGP offers the most reliable point predictions, evidenced by its lowest RMSP. The HGP also maintains nominal coverage for its prediction intervals. While the AGP2 achieves similar coverage, its interval scores are considerably higher. Table S.6 in the Supplementary Material provides a breakdown by data type, highlighting the poor performance of competing methods for point-referenced data. The PI frequentist coverage, for example, can be as low as 3%.
In Table S.7 of the Supplementary Material, we further compared the predictive performance of those models to two additional alternatives: a Log-Normal model with HGP random effects, and a GP where the areal units are reduced to their centroids, resulting in a geostatistical model. We have not tested an AGP where the response follows a Log-Normal distribution, because the random effects would need to be aggregated on the response scale (suen2025coherent), which can be seen as a consequence of Jensen’s inequality. Surprisingly, the AGP models performed the worst, while the Log-Normal HGP provided the second-best predictive performance.
The results from the AGP models diverged significantly depending on the mesh resolution. The credible interval width for the intercept varies significantly between the two models. The AGP1 model indicates moderate spatial dependence, while AGP2 suggests weak dependence. This weak dependence may imply AGP2 assigns nearly independent random intercepts to each data point, with variances inversely proportional to the sample unit’s area. AGP mesh resolution is typically practitioner-informed, and results are highly sensitive to this tuning parameter, which lacks a consensus on “optimal resolution” (fuentes2005model; liu2011empirical; wang2021combining).
@@ 6 Discussion
The proposed framework unifies the modeling of spatial data, extending geostatistical models to include areal and fused data within the same setting. Despite the mathematical simplicity, the subtle change of paradigm powered by the HGP opens a new direction of research in spatial statistics while taking advantage of what has been proposed in the geostatistics and GP literature. The proposal has the potential to simplify how we deal with complex problems such as spatial misalignment and change of support. In simpler scenarios, as a prior for spatial random effects in a GLMM, the HGP proves to be competitive to models specially designed for areal and fused data types. Our simulation studies indicate that the HGP is particularly advantageous for areal data with units of varying sizes and shapes, as confirmed by our disease mapping application. Unlike the traditional areal models, the HGP also provides a directly interpretable spatial parameter, offering insights into the practical range of spatial dependence. In data fusion contexts, the HGP delivers more reliable interval predictions compared to the sparse AGP, and it provided similar results to the fine AGP in parameter estimation and spatial range. Additionally, in the air pollution data application, the proposed model consistently outperformed the AGP regardless of the resolution of its mesh. This success is even more remarkable considering the HGP bypasses the subjective step of defining a grid or mesh, a crucial but problematic requirement for the AGP.
The empirical assessment in Section 2.3 reveals an interesting insight: the properties of the induced process appear linked to the geometric complexity of the sets involved. This is most evident in the contrasting results from our two scenarios. For the data fusion scenario, which involves only convex sets, all correlation matrices were positive definite. In contrast, for the areal data scenario, where most spatial units are non-convex, we observed negative eigenvalues at the border of the parameter space (Figure 1). This contrast provides strong empirical evidence that the PEC function is a valid and robust choice for applications involving convex sets, such as most data fusion problems.
While our extensive empirical assessments consistently demonstrate the positive definiteness of the PEC function across practical parameter ranges, a formal proof establishing the general conditions under which functions of the Hausdorff distance are positive definite remains an open and challenging theoretical question. Nevertheless, the proposed HGP addresses several long-standing problems in spatial statistics. For instance, traditional areal models are unsuitable for out-of-sample prediction and often have counter-intuitive parameters interpretation (wall2004close; assuncao2009neighborhood). Furthermore, the validity of CAR models, for instance, also depends on constraining the parametric space to ensure a non-singular precision matrix (wall2004close). While data fusion models do not share the same issues, we have shown that their results can be significantly affected by the arbitrary precision of required approximations. By introducing the HGP and demonstrating its utility, we aim to provide a foundation for future theoretical work, thereby bridging the gap between practical application and complete mathematical understanding.
Several conceptual challenges for the HGP remain open for future investigation. Relaxing the isotropy assumption, for instance, requires further theoretical development. To apply the HGP to data on manifolds like the sphere, new results are needed to guarantee the process’s validity. Furthermore, because the HGP is a set-valued function, any analysis of its curvature (e.g., halder2024Bayesian) necessitates more general concepts of differentiation (khastan2021new). It is important to note, however, that this latter complexity does not prevent the use of powerful inference frameworks that rely on the posterior’s gradient, such as variational inference (blei2017variational).
Although the $\mathcal{O}(n^{3})$ computational complexity is a limitation, the method was developed with the intent of leveraging what has already been developed for GPs. Thus, adapting scalable nearest neighbor approximations (datta2016hierarchical; quiroz2023fast) for the HGP is a viable avenue for future research. The main challenge would be to identify nearest neighbors in the Hausdorff distance sense while avoiding the storage of a $n\times n$ distance matrix.
Conflict of interest
All authors declare that they have no conflicts of interest.
Appendix A Review of Existing Methods
The BYM model we use corresponds to a zero-mean multivariate Normal model with covariance matrix defined as
$\sigma^{-2}((1-\xi)\bm{I}+\xi\bm{Q}^{-}),$
where $\sigma$ is the average marginal SD, $\bm{I}$ is the identity matrix, and $\bm{Q}^{-}$ is the generalized inverse of the scaled precision matrix of an ICAR model. The spatial dependence parameter $\xi$ is a mixing parameter. Its parametric space comprises of real numbers between 0 and 1, with larger values indicate stronger spatial dependence. The entries of $\bm{Q}$ are as follows:
$Q_{ij}=\begin{cases}kn_{N(i)},&\text{ if }i=j,\\ -k,&\text{ if }i\sim j,\\ 0,&\text{ elsewhere,}\end{cases}$
where $n_{N(i)}$ denotes the number of neighbors for the region $i$, and $i\sim j$ indicates that $j$ and $i$ are neighbors, and $k$ is the scaling constant. The scaling constant depends on the adjacency matrix and enhances the interpretation of $\sigma$, which becomes the average marginal standard deviation, as opposed to the conditional standard deviation in the original formulation (riebler2016intuitive). We used the BYM as a prior for the random effects in the spatial GLMMs (Section 3.1) for areal data in Section 5.1. We also generated data from a spatial GLMM assuming the BYM was the actual distribution of the spatial random effects for part of the simulation studies in Section 4.1.
The DAGAR model, in our context, is a zero-mean multivariate Normal model. We used the DAGAR as a prior for the spatial GLMM (Section 3.1) random effects (e.g., $\mathbf{z}$) for areal data, in particular, in Section 4.1 and 5.1. The precision matrix for the DAGAR model is constructed using the data ordering, an adjacency matrix, and two parameters: a conditional precision, $\tau$, and a spatial dependence parameter, $\psi$. The $\psi$ parameter represents the average spatial correlation between adjacent spatial units. Specifically, the data ordering and adjacency matrix define a tree graph, and the precision matrix is derived from an order-1 autoregressive process on the nodes of this graph. A key advantage of this model is its computational efficiency, as it can be evaluated using either closed-form conditional distributions or direct formulas for the Cholesky decomposition of its covariance matrix. In particular, the Cholesky decomposition of the precision matrix under a DAGAR model can be written as $\mathbf{F}^{-1/2}(\mathbf{I}-\mathbf{B})$, where $\mathbf{F}$ is a diagonal matrix, $\mathbf{I}$ is the identify matrix, and $\mathbf{B}$ is lower-triangular (datta2022nearest). The entries of $\mathbf{F}$ and $\mathbf{B}$, respectively, are given by
$F_{ii}$ $={[\tau(1+1+n_{\vec{N}(i)}\psi^{2})]}^{-1}\quad\text{ and }$
$B_{ij}$ $=\mathbbm{1}(j\sim i,j<i)\frac{\psi}{1+1+n_{\vec{N}(i)}\psi^{2}},$
where $\mathbbm{1}(A)$ denotes an indicator function, equaling $1$ if the event $A$ happens and 0 otherwise. Note that, in this case, $n_{\vec{N}(i)}=\sum_{j}\mathbbm{1}(j\sim i,j<i)$ represents the number of “directional neighbors”. Although the DAGAR depends on an arbitrary order for the data points, the authors also developed an “order-free” variant of the model. However, the “order-free” DAGAR does not enjoy the computational efficiency of its ordered counterpart.
The AGP model is employed as a prior on the random effects ($\mathbf{z}$) of the spatial GLMM (Section 3.1) for fused data. In particular, we used this model to generate data for the simulation studies in Section 4, and to analyze the dataset from Section 5.2. The AGP core assumption is as follows:
[…]
where $\{Z(\mathbf{s})\,:\,\mathbf{s}\in D\}$ is a (point-level) zero-mean, second order stationary and isotropic GP (cressie1993statistics, p. 53) defined over the study region, $\mathcal{A}(\cdot)$ is a function that returns the area of a spatial object.